How do you write an equation for an upside down parabola going through points (0,0) (24,12) and (32,0)?
Answer by lwsshak3(11628)

How do you write an equation for an upside down parabola going through points (0,0) (24,12) and (32,0)?
***
Set up a system of 3-equations using form y=Ax^2+Bx+C
(0,0) 0=0+0+C
C=0
(24,12) 12=24^2A+24B+0
(32,0) 0=32^2A+32B+O
..
384=32*24^2A+768B+0
0=24*32^2A+768B+0
..
384=18432A+768B+0
0=24576A+768B+0
subtract:
384=-6144A
A=-6144/384=-1/16
32B=-32^2A=-32^2*-1/16=64
B=2
Equation:
y=-x^2/16+2x
complete the square:
y=(-1/16)(x^2-32+256)+16
y=(-1/16)(x-16)^2+16
This is an equation of a parabola of the form: y=A(x-h)^2+k, (h,k)=(x,y) coordinates of center.
For given parabola:
vertex: (16,16)
A=-1/16